Circular No. 8360

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004ci AND 2004cr R. J. Foley, M. Ganeshalingam, and A. V. Filippenko, University of

California, report that CCD spectra (range 330–1000 nm), obtained on June 20 UT with the Shane 3-m telescope at Lick Observatory, show that SNe 2004ci (IAUC 8357) and 2004cr (IAUC 8359) are both of type II. The expansion velocities, derived from the minimum of Hβ for SNe 2004ci (recession velocity 4158 km/s from narrow Hα) and 2004cr (recession velocity 5055 km/s from narrow Hα), are 11813 and 8443 km/s, respectively.

T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 340–730 nm) of SN 2004ci, obtained by E. Adams and M. Westover on June 19.23 UT with the Mt. Hopkins 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova. The spectrum consists of a blue continuum with P-Cyg lines of hydrogen and helium. Adopting the NASA/IPAC Extragalactic Database recession velocity of 4092 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 10700 km/s.

COMET C/2001 Q4 (NEAT) Further to IAUC 8358, R. W. Russell, D. L. Kim, M. L. Sitko, and

W. J. Carpenter, report additional 3–13-μm spectrophotometry of C/2001 Q4, obtained on June 20.2 UT at Mt. Lemmon (integration time 45 min; reference star α Boo): “A smooth comet continuum was seen to rise from 3.5 to 8.4 μm, beyond which a moderate silicate emission band was observed. An underlying blackbody continuum with a temperature of ∼ 315 ± 5 K was fit to the continuum flux at 5, 8.4, and 12 μm. This grain temperature is ∼ 21 ± 2 percent higher than that of an equilibrium blackbody at the heliocentric distance of the comet. The silicate-feature-to-continuum ratio was 1.32 ± 0.02 — slightly greater than the value of 1.26 reported on June 17.2 (IAUC 8358), but still significantly less than that reported on May 31.2 (1.43; IAUC 8351) and May 14.2 (1.65; IAUC 8342). The crystalline olivine feature at 11.2 μm continues to be present with approximately the same contrast as originally seen. With our aperture, the comet has the following narrowband (∼ 0.25 μm) magnitudes and combined random errors: [3.7 μm] = 7.73 ± 0.17, [4.7 μm] = 5.49 ± 0.20, [5 μm] = 4.95 ± 0.23, [8 μm] = 2.10 ± 0.18, [10.5 μm] = 0.50 ± 0.05, [12 μm] = 0.26 ± 0.04.”

Visual total-magnitude estimates: June 8.89 UT, 6.6 (W. Hasubick, Buchloe, Germany, 10×50 binoc.); 19.93, 6.4 (A. Diepvens, Balen, Belgium, 20×50 binoc.); 21.92, 6.7 (J. J. Gonzalez, Asturias, Spain, 7×50 binoc.).

2004 June 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green